# CHIME-o-Grav: Wideband Timing of Four Millisecond Pulsars from the NANOGrav 15-yr dataset

### v2

@@ 5.2 CHIME Frequency coverage and Wideband TOA precision
The corresponding WB DMs are the opposite, with the CHIME WB DMs being much more precise than GUPPI or GASP DMs. This decrease is likely due to CHIME’s frequency coverage being larger and covering lower frequencies than any of the GBT receivers used. DM is measured by the delay between pulse arrivals from different frequencies. Lower frequencies will have a much larger lever arm on measuring DM delay, which scales with $\nu^{-2}$. Using PSR J0645++5158 as an example, the DM delay across the CHIME bandwidth was 291 ms, whereas the delay across the 820 MHz receiver was 43 ms and the delay across the L-Band receiver would only be 29 ms.
@@ 6 Conclusions
The inclusion of CHIME data in PTA datasets presents a unique laboratory for studying the effect of the solar wind on pulsar timing. The more precise DM measurements we get from the low-frequency, high cadence CHIME data are extremely useful for illustrating short time scale ISM fluctuations, particularly DM variabilty from the solar wind. This was particularly illustrated in the DM data for PSR J2145-0750. In future analyses CHIME data could be used to improve global solar wind models.
@@ Appendix A The chimera pipeline
We developed the chimera pipeline [^6] Available at https://github.com/abhisrkckl/chimera. to measure wideband TOAs from the partially folded profiles in the Timer format generated by the CHIME/Pulsar backend. chimera is written in Python and provides a command-line script named chimerawb777chimera is written with a structure similar to and reuses code from the pinta pipeline of the Indian Pulsar Timing Array (Susobhanan et al., 2021). . The chimerawb command takes an input directory containing Timer archives (van Straten & Bailes, 2011), a configuration file in the JSON format, optionally a meta file containing the list of input archives to be processed, and an output directory as input, and writes a tim file containing wideband TOAs in the IPTA format into the output directory. The configuration file contains information about the pulsar name, fiducial DM, template portrait, number of subintegrations/channels for time/frequency scrunching, and a list of bad channels for pulsar-specific RFI flagging. The template portrait is generated manually beforehand using the ppspline command of PulsePortraiture from an epoch-averaged profile obtained from multiple observations, since individual CHIME observations may have low S/N (Pennucci, 2019); see section 2.4. The TOAs are generated using the pptoa command of PulsePortraiture (Pennucci, 2019).

### v3

@@ V.2 CHIME Frequency coverage and WB TOA precision
The corresponding WB DMs are the opposite, with the CHIME WB DMs being much more precise than GUPPI or GASP DMs. This decrease is likely due to CHIME’s frequency coverage being larger and covering lower frequencies than any of the GBT receivers used. DM is measured by the delay between pulse arrivals from different frequencies. Lower frequencies will have a much larger lever arm on measuring DM delay, which scales with $\nu^{-2}$. Using PSR J0645++5158 as an example, the DM delay across the CHIME bandwidth was 291 ms, whereas the delay across the 820 MHz receiver was 43 ms and the delay across the L-Band receiver would only be 29 ms. ISM propagation effects are a significant source of noise for PTA analyses, so the increase in total observing range due to the inclusion of low-frequency, high precision, CHIME DM data will be important for modeling frequency-dependent DMs (Agazie et al., 2023c). The high cadence observing also is beneficial for monitoring short and long timescale DM fluctuations.
@@ VI Conclusions
As PTAs become more sensitive, with both longer timing datasets and increased observing bandwidths, the potential for future detection of continuous waves from individual supermassive black hole binaries will grow. The inclusion of CHIME data in PTA datasets is particularly important for studying interstellar propagation effects which can significantly impact PTA sensitivity. In future NANOGrav datasets, CHIME data will help mitigate the loss of low-frequency observing due to the collapse of the Arecibo Telescope. CHIME WB DM data also presents a unique laboratory for studying the effect of the solar wind on pulsar timing. The more precise WB DM measurements we get from the low-frequency, high cadence CHIME data are extremely useful for illustrating short time scale ISM fluctuations, particularly DM variability from the solar wind. This was particularly well-demonstrated in the DM data for PSR J2145-0750. In future analyses CHIME data could be used to build more precise global solar wind models.
@@ Appendix A The chimera pipeline
We developed the chimera pipeline [^6] Available at https://doi.org/10.5281/zenodo.18458818. to measure wideband TOAs from the partially folded profiles in the Timer format generated by the CHIME/Pulsar backend (Susobhanan, 2026). chimera is written in Python and provides a command-line script named chimerawb777chimera is written with a structure similar to and reuses code from the pinta pipeline of the Indian Pulsar Timing Array (Susobhanan et al., 2021). . The chimerawb command takes an input directory containing Timer archives (van Straten & Bailes, 2011), a configuration file in the JSON format, optionally a meta file containing the list of input archives to be processed, and an output directory as input, and writes a tim file containing wideband TOAs in the IPTA format into the output directory. The configuration file contains information about the pulsar name, fiducial DM, template portrait, number of subintegrations/channels for time/frequency scrunching, and a list of bad channels for pulsar-specific RFI flagging. The template portrait is generated manually beforehand using the ppspline command of PulsePortraiture from an epoch-averaged profile obtained from multiple observations, since individual CHIME observations may have low S/N (Pennucci, 2019); see section II.4. The TOAs are generated using the pptoa command of PulsePortraiture (Pennucci, 2019).